OpenStudy (anonymous):

if f(x)=3x^2-2x-5, find(2u+1) can you explain how to work the problem and whats the answer i can get to f(2u+1)=3(2u+1) after that i don't know

OpenStudy (radar):

\[f(2u+1)=3(2u+1)^{2}-2(2u+1)-5\] Thats a start.

OpenStudy (radar):

\[=3(4u ^{2}+4u+1)-4u-2-5\] just keep on trucking!

OpenStudy (radar):

Earth to shrryflores, do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Let's continue then: \[12u ^{2}+12u+3-4u-7\] \[12u ^{2}+8u+4\]

OpenStudy (radar):

change that +4 to -4

OpenStudy (anonymous):

ok it makes since

OpenStudy (radar):

since there was no = sign I think that is as far as needed\[12u ^{2}+12u-4\]

OpenStudy (radar):

\[f(2u+1)=12u ^{2}+12u-4\]

OpenStudy (anonymous):

\[f(2u+1)=12u^2+8u-4\] You can find the roots of u using the quadratic formula: \[u=-\frac{1}{3}\pm \frac{2}{3}\rightarrow u=-1, u=\frac{1}{3}\]

OpenStudy (radar):

I was wondering if it was to be taken any further? Couldn't determine if they were wanting the roots.

OpenStudy (anonymous):

I gave the roots just in case

OpenStudy (radar):

Fine business, they say "whats the answer" maybe that was what they wanted. thanks and good luck to you and shrryflores